OpenStudy (anonymous):

The endpoints of AB⎯⎯⎯⎯⎯ are A(1, 4) and B(5, -1). What are the coordinates of the point that divides AB⎯⎯⎯⎯⎯ in the ratio 2 : 3?

OpenStudy (campbell_st):

this is the way I do the question \[\frac{2}{3} = \frac{A}{B}\] cross multiply 2B = 3A so then is 2(5, -1) = 3(1, 4) so for x \[x = \frac{2 \times 5 + 3 \times 1}{2 + 3}\] and \[y = \frac{2 \times -1+ 3 \times 4}{2 + 3}\] now just simplify them...